OpenStudy (anonymous):

Please Help Me 45/9 + m X 3 = 23

OpenStudy (gw2011):

45/9+3m=23 5+3m=23 3m=23-5 3m=18 m=6

OpenStudy (anonymous):

Thank-you soo much

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

3 X [5 + (7-5) exponent 2]=

OpenStudy (gw2011):

3[5+(7-5)^2] 3[5+(2)^2] 3(5+4) 3(9)=27